CSWSlearning:

A school puts on a play. The play costs $1,200 in expenses. The students charge $4.00 for tickets. There will be one performance of the play in an auditorium that seats 500 people. What is the domain of the function that shows the profit as a function of the number of tickets sold?

Ultrilliam:

@Zowie

CSWSlearning:

Also this one Your friend has $90 when he goes to the fair. He spends $5 to enter the fair and $15 on food. Rides at the fair cost $1.50 per ride. Which function can be used to determine how much money he has left over after x rides?

Zowie:

Can you type/insert the answer choices?

CSWSlearning:

yes I may sorry, First question answers... A.) The domain is all real numbers from negative 1,200 to positive 2,000 B.) The domain is all integers from negative 1,200 to positive 2,000. C.) The domain is the integers from 0 to 500. D.) The domain is all real numbers from 0 to 500. Second question answers... A.) f(x) = 1.5x + 70 B.) f(x) = −1.5x + 70 C.) f(x) = −1.5x − 70 D.) f(x) = −20x + $90

CSWSlearning:

Thanks for the help! :)

Zowie:

The first one I believe is D. The second one is definitely C.

CSWSlearning:

Could you explain your reasoning for question #1

Zowie:

Honestly it's what I think. Not a 100% correct answer. But the second one is definitely C. I know that for a fact.

CSWSlearning:

Ok...

CSWSlearning:

The answer for number 1 was C and the second one was B :) thanks for trying though

dude:

Question 1 could've been C because you can only accommodate 500 people and you cant have a fraction of a person, therefore so not ALL real numbers would fit